I need help! This is due tomorrow! (It's 7th grade math by the way.)

The question says: Write an equation to find the value of x so that each pair of polygons has the same perimeter. Then solve.
1. A triangle has sides that say: x+6, x+9, and x+3. A rectangle beside it has sides that say x+4 and x+1.
2. A hexagon has sides that all read 5x. A triangle next to it has sides that say x+14, x+17, and 8x+9.
PLEASE help me!

Answer by Earlsdon(6294)

1) For the triangle, let the perimeter = P, then...
(x+6)+(x+9)(+x+3) = P Simplify.
3x+18 = P
For the rectangle, let the perimeter = P (the same P), then...
2(x+4)+2(x+1) = P Simplify.
2x+8+2x+2 = P
4x+10 = P Now, because P = P you set these two equations equal and solve for x...
3x+18 = 4x+10 Subtract 3x from both sides.
18 = x+10 Now subtract 10 from both sides.
8 = x or x = 8.
Now you can apply this to the next problem.